OpenStudy (anonymous):

The answer is i, right? ;) Simplify i 30 1 i -i -1

OpenStudy (anonymous):

is 30 exponential?

OpenStudy (anonymous):

if so, then it all simplifies to just -1

OpenStudy (anonymous):

Yes, my error

OpenStudy (anonymous):

just like i^2 which will be -1

jimthompson5910 (jim_thompson5910):

30/4 = 7 remainder 2 since you have a remainder of 2, i^(30) = i^2 since i^2 = -1, i^(30) = -1

OpenStudy (anonymous):

That makes SO MUCH sense, thanks! :)

OpenStudy (anonymous):

I've gotten that confused several times